OpenStudy (anonymous):

lim as dx--->0. (5(x+dx)-(x+dx)-(5x^2-x))/dx

OpenStudy (pradipgr817):

how about infinity

OpenStudy (anonymous):

simplify the numerator 1st...

OpenStudy (anonymous):

well, i'd like to know the process, you know? like the distributing & cancelling out? Because i got to 5x^3+15x^2dx-15xdx^2+5dx^3-dx-5x^2 and idk if that's right or what to do next.

OpenStudy (anonymous):

@akash123 is that simplification right? ^^^

OpenStudy (anonymous):

let's do the simplification together..

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

1st two terms of the numerator...5(x+dx)-(x+dx)

OpenStudy (anonymous):

5x+5 dx - x-dx does it make sense?

OpenStudy (anonymous):

5(x+dx)-(x+dx) = 5 * (x+dx)-1 * (x+dx)

OpenStudy (anonymous):

i forgot to add that the (x+dx) is square.d, sorry. but yes that makes sense.

OpenStudy (anonymous):

the first one.

OpenStudy (anonymous):

ok ...write the complete question again..

OpenStudy (anonymous):

(5(x+dx)^2 -(x+dx)-(5x^2-x))/dx

OpenStudy (anonymous):

now it's ok?

OpenStudy (anonymous):

(5(x+dx)^2 - (x+dx) - (5x^2-x))/dx

OpenStudy (anonymous):

you know the identity (a+b)^2 = a^2 + b^2 + 2 a b

OpenStudy (anonymous):

yessss sir

OpenStudy (anonymous):

so using this identity 1st term of the numerator 5(x+dx)^2 = 5( x^2+ dx ^2 +2 x*dx)

OpenStudy (anonymous):

I m not a sir...i m akash only

OpenStudy (anonymous):

sorry I say that, i wasn't calling you a sir. and yes, I'm following you.

OpenStudy (anonymous):

5(x+dx)^2 -(x+dx)-(5x^2-x) = 5( x^2+ dx ^2 +2 x*dx) - x -dx - 5 x^2 +x

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

5( x^2+ dx ^2 +2 x*dx) - x -dx - 5 x^2 +x = 5x^2+ 5dx ^2 + 10 x*dx- x -dx - 5 x^2 +x

OpenStudy (anonymous):

now look for similar terms..

OpenStudy (anonymous):

ohhhhhhh 5x^2 & Xs

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

5dx ^2 + 10 x*dx-dx

OpenStudy (anonymous):

it's ok?

OpenStudy (anonymous):

5x^2 - 5x^2+x -x ...got cancelled

OpenStudy (anonymous):

yes, I realize the simplified version i posted corresponded to another problem, yes I got that. what's next?

OpenStudy (anonymous):

so u have 5dx ^2 + 10 x*dx-dx in the numerator after simplification

OpenStudy (anonymous):

now take dx common from numerator...

OpenStudy (anonymous):

i was not underestimating u...if u think so...sorry

OpenStudy (anonymous):

dx ( 5dx + 10 x-1)

OpenStudy (anonymous):

5dx(x+2x-1)/ dx so you cancel the dx on top & bottom then you plug in? it's fine, i didn't think that

OpenStudy (anonymous):

ohhhh i see what you did.

OpenStudy (anonymous):

after taking dx common u'll get dx ( 5dx + 10 x-1)

OpenStudy (anonymous):

yeah why don't you take out the 5?

OpenStudy (anonymous):

u can take also 5 common...u'll be left with dx 5( dx + 2 x-1/5)

OpenStudy (anonymous):

ok so same thing?

OpenStudy (anonymous):

nevermind. what's next?

OpenStudy (anonymous):

cancel the dx from numerator and denominator...

OpenStudy (anonymous):

and put dx=0

OpenStudy (anonymous):

after canceling ( 5dx + 10 x-1) or 5( dx + 2 x-1/5)

OpenStudy (anonymous):

one is with dx common and other is by taking common 5 dx

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

now plug in dx=0

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

no...10x-1

OpenStudy (anonymous):

u have 10 x also in the expression...

OpenStudy (anonymous):

ohhh ok. so what do you do with the 10x-1?

OpenStudy (anonymous):

it's the final answer of your question

OpenStudy (anonymous):

as u have been asked....lim as dx--->0

OpenStudy (anonymous):

oh ok so you know you're at the final answer when there are no more dxs?

OpenStudy (anonymous):

yes...u have been given an expression ...(5(x+dx)^2 - (x+dx) - (5x^2-x))/dx...and u are being asked...what will be value of the expression if dx--->0...so there should be no dx in ur final answer.

OpenStudy (anonymous):

u can try this limit lim x---->2 (x^2-4)/(x-2)

OpenStudy (anonymous):

lim x---->2 (x^2-4)/(x-2)

OpenStudy (anonymous):

lim=4

OpenStudy (anonymous):

yes...perfectly right...as u r seeing here there is no x in ur final answer.

OpenStudy (anonymous):

similarly there wont be any dx in the answer of your question...and x will be treated as a constant..

OpenStudy (anonymous):

oh ok. do you think you can help me with one more question?

OpenStudy (anonymous):

yes...ask

OpenStudy (anonymous):

sorry my computer went crazy on me

OpenStudy (anonymous):

ok so it's 5x^3-10x^2+12

OpenStudy (anonymous):

and I know you have to plug in x+dx in the xs

OpenStudy (anonymous):

and I know you have to plug in x+dx in the xs didnt get it?

OpenStudy (anonymous):

like 5(x+dx)^3...

OpenStudy (anonymous):

got it...u have to plug (x+dx) in the place of x everywhere and plug in dx=0..is it so?

OpenStudy (anonymous):

i don't know about the dx =0 but after distributing the 5 and 10 I got

OpenStudy (anonymous):

write down the complete question...

OpenStudy (anonymous):

find lim as dx----> 0. (f(x +dx) - f(x))/dx. those are the directions.

OpenStudy (anonymous):

actually this id definition of f'(x)

OpenStudy (anonymous):

then the problem is 5x^3-10x^2-12

OpenStudy (anonymous):

and u have to calculate f'(x) for a given f(x)

OpenStudy (anonymous):

now i got the question..

OpenStudy (anonymous):

here f(x) = 5x^3-10x^2-12

OpenStudy (anonymous):

and u have to find f'(x) using f'(x)= dx----> 0. (f(x +dx) - f(x))/dx

OpenStudy (anonymous):

now 1st calculate f(x+dx)=5(x+dx)^3-10(x+dx)^2-12

OpenStudy (anonymous):

yeah I got up to 15x^2+15xdx+25dx^2-20x+10dx, but idk if i did it right, or made mistakes along the way. or what to do next.

OpenStudy (anonymous):

no...there must be some cubic term in ur expression after simplification

OpenStudy (anonymous):

let's do the calculation together...there are two ways

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5(x+dx)^3-10(x+dx)^2-12 = 5(x+dx)^2 *{ (x+dx)-2} -12

OpenStudy (anonymous):

i have taken 5(x+dx)^2 common from 1st two terms...

OpenStudy (anonymous):

it's ok?

OpenStudy (anonymous):

seems reasonable...

OpenStudy (anonymous):

x+dx-1?

OpenStudy (anonymous):

how did u get it?

OpenStudy (anonymous):

i mean 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok that makes it a lot simpler

OpenStudy (anonymous):

5(x+dx)^2 *{ (x+dx)-2} -12 =5( x^2 + dx^2 +2 x dx)*( x+dx -2) -12

OpenStudy (anonymous):

now using distributive law do the multiplication..

OpenStudy (anonymous):

5 x^2 +5 dx^2 +10 x dx)*( x+dx -2) -12

OpenStudy (anonymous):

is the -12 diistributed as well?

OpenStudy (anonymous):

5( x^2 + dx^2 +2 x dx)*( x+dx -2) -12 5{ x^2*( x+dx -2) + dx^2* ( x+dx -2) + 2 x dx*( x+dx -2) } -12

OpenStudy (anonymous):

oh you distribute the 5 through all that?

OpenStudy (anonymous):

no...i didnt distribute 5 yet...it's still outside..